# WinoPron: Revisiting English Winogender Schemas for Consistency, Coverage, and Grammatical Case

### v2

@@ 5 Measuring Pronominal Bias
We apply our method to SpanBERT to evaluate bias with he/him/his, she/her/her, they/them/their, and xe/xem/xyr. Table 3 shows results for SpanBERT-large, where we find that pronominal bias does not always match across grammatical case, i.e., positive bias with xe for an occupation does not entail positive bias with xyr. In fact, even contradictory associations are possible; SpanBERT-base has a positive bias between manager and them, but a negative bias betweeen manager and their. Across pronoun series, we find that pronominal bias is non-exclusive, i.e., SpanBERT-large has positive bias between baker and her, their as well as xyr, and additionally, these results do not imply that the model must have a negative bias between baker and his. This further highlights the need for evaluation that goes beyond binary, oppositional operationalizations of gender via pronouns.
@@ 6 Conclusion
We demonstrate a number of issues with the well-known Winogender schemas dataset, which we fix in our new, expanded WinoPron dataset. In addition, we propose a novel way to evaluate pronominal bias in coreference resolution that goes beyond the binary and focuses on lexical gender. With our new dataset and method, we evaluate both supervised coreference resolution systems and language models, and find that our changes impact performance and bias characteristics. Our work demonstrates that measurements of bias and robustness are only as good as the datasets and metrics we use to measure them, and we call for careful attention when developing resources for and discussing performance on coreference resolution.
@@
As in Winogender schemas, our schemas are not “Google-proof” and could conceivably be solved with heuristics, including word co-occurrences, which is a primary concern when creating and evaluating Winograd schemas (Levesque et al., 2012; Amsili and Seminck, 2017; Elazar et al., 2021). The fact that we do not control for this means that our dataset gives generous estimates of model performance and bias, but it also means that this dataset is inappropriate to test model “reasoning.” Our dataset construction instead controls for simple system heuristics that are relevant for coreference resolution, such as always picking the first entity in the sentence, or always picking the second. In interpreting system performance, we also echo Rudinger et al.’s (2018) view of Winogender schemas as having “high positive predictive value and low negative predictive value.” In other words, that they may demonstrate evidence of pronominal bias in systems, but not prove its absence.
We take steps to prevent data contamination following Jacovi et al. (2023), including not releasing our data in plain text, and not evaluating with models behind closed APIs that do not guarantee that our data will not be used to train future models (Balloccu et al., 2024). However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on WinoPron with models trained on data after August 2024.
Finally, we note that as our evaluation set only contains one set of templates per occupation-participant pair, our results represent a point in the distribution of bias related to that occupation. In the case of large language models in particular, using a small number of templates for templatic evaluation is known to be brittle even to small, meaning-preserving changes to the template (Seshadri et al., 2022; Selvam et al., 2023). Our dataset’s small size is a result of us requiring a tightly controlled and structured dataset to evaluate how coreference resolution varies. Thus, it may differ from realistic examples (which would have other differences that confound bias results). We wish to emphasize that in addition to controlled datasets like ours, realistic evaluation is also necessary for holistically evaluating performance, robustness and bias in coreference resolution.
@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, and are grateful to Rachel Rudinger and Benjamin Van Durme for their comments. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Anne Lauscher’s work is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A Related work
Appendix B List of Occupations
@@
Appendix C Annotator Demographics
@@
Appendix D Annotation Instructions
D.1 Task 1 Description
@@ D.1.1 Grammaticality
D.1.2 Questions about pronouns
@@ D.2 Task 2 Description
Appendix E Prompting
  1. 1.
    No options
  2. 2.
    The occupation is presented first and the participant second
  3. 3.
    The participant is presented first and the occupation second
@@ Appendix E Prompting
Table 4: Prompting templates, where “task” is filled with each dataset instance, “pronoun” is the only third person singular pronoun that appears in that dataset instance, and “options” is a list of two options to choose from–the occupation and the participant.
Appendix F Data Creation and Validation Details
Constructing the schemas was a challenging and time-consuming process as we ran into intermittent issues, primarily relating to ambiguity (§F.1) and verbal constraints (§F.2), which we had to iteratively fix. In addition to these types of errors, which could only be detected and fixed by humans, we also had structural constraints that could be programmatically checked. We wrote checks for these constraints in Python and describe them in §F.3.
F.1 Ambiguity
One of the big difficulties we faced in creating our schemas was ensuring that we created unambiguous instances of singular specific they. For example, if an advisor and student were meeting to discuss their future, this could also potentially refer to their future together. This problem applied to all grammatical cases (nominative, accusative, and possessive). In addition, possessive sentences were potentially ambiguous across all pronoun series; when discussing a doctor and a patient and someone’s diagnosis, this could be the doctor’s diagnosis (i.e., the diagnosis that was made by the doctor), or the patient’s diagnosis (i.e., the diagnosis that the patient received.
F.2 Verbal Constraints
In some cases the requirement that sentences should be structurally identical until the pronoun created difficulties in finding appropriate, logically and semantically correct endings for the two sentences. It was sometimes difficult to create single-entities sentences that were semantically close to the two-entity versions because the original situation only made sense with two entities (e.g. sentences with "X told Y"). A smaller issue was to ensure that the verbs were congruent with all nominative pronouns. Therefore, we had to only used past tense and avoid the use of was/were.
F.3 Automatic Checks
We automatically checked our data for completeness, i.e., that every occupation-participant pair had sentence templates for nominative, accusative, and possessive grammatical cases. We also checked structural constraints, e.g., that a pair of templates must always be identical until the pronoun slot. We also ensured that there was no ambiguity introduced through additional pronouns appearing in the sentence, limited to the set of pronouns we are interested in, i.e., not including it/its pronouns. Using the human evaluation data, we checked that in every template, the pronoun corefers with what we want it to corefer, according to our data label.
Appendix G Additional Results
Performance on neopronouns (xe/xem/xyr) is markedly worse compared to other pronouns for SpanBERT, CAW-coref and LingMess, but not for the larger sizes of FLAN-T5, as shown in Table 5.
Pronouns LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
he/him/his 0.79 0.80 0.63 0.71 0.52 0.64 0.81 0.89 0.90
she/her/her 0.76 0.77 0.62 0.72 0.52 0.66 0.83 0.91 0.90
they/them/their 0.68 0.64 0.61 0.69 0.52 0.61 0.77 0.89 0.88
xe/xem/xyr 0.09 0.39 0.60 0.69 0.51 0.59 0.70 0.88 0.87
Table 5: F1 scores of coreference resolution systems on WinoPron, split by pronoun series. Red, italicized numbers are worse than chance.
Table 6 shows how coreference resolution systems perform on the single-entity templates in our dataset, which are meant to be a very simple baseline. Here, SpanBERT-base emerges as a clear winner, with near-perfect performance across all cases and pronoun sets.
@@ Appendix G Additional Results
All 0.73 0.76 0.96 0.88 0.77 0.76 0.81 0.83 0.84
Nominative 0.80 0.82 0.99 0.99 0.78 0.90 0.85 0.93 0.97
Accusative 0.61 0.65 0.87 0.68 0.76 0.70 0.70 0.70 0.67
Possessive 0.77 0.78 1.00 0.97 0.78 0.78 0.85 0.86 0.89
he/him/his 0.93 0.94 0.95 0.86 0.78 0.81 0.87 0.88 0.89
she/her/her 0.91 0.92 0.96 0.89 0.77 0.81 0.88 0.89 0.87
they/them/their 0.75 0.70 0.96 0.89 0.79 0.76 0.84 0.86 0.87
xe/xem/xyr 0.02 0.27 0.95 0.89 0.75 0.66 0.67 0.69 0.75
Table 6: F1 scores of coreference resolution systems on single-entity sentences in WinoPron. In addition to total F1, we report performance split by grammatical case and by pronoun series.

### v3

@@ 5.2 Results
Even though SpanBERT-large has positive bias for baker and her, their as well as xyr, this does not imply that the model must have a negative bias between baker and his; it does not. This further highlights the need for evaluation that goes beyond binary, oppositional operationalizations of gender via pronouns.
6 Discussion
By systematically identifying and fixing issues with Winogender Schemas (Rudinger et al., 2018), we create a new dataset, WinoPron, and find that: (1) different grammatical cases of pronouns show vastly different performance and bias characteristics, (2) pronominal biases are rich and varied, of which he and she are only the tip of the iceberg, and (3) model biases are complex and do not necessarily match our intuitions about them. Based on our findings, we make some recommendations for researchers who study coreference resolution and those who study bias and fairness via pronouns.
First, grammatical case is a dimension of pronominal performance and bias that warrants more study (Munro and Morrison, 2020). In particular, we hope that future work further investigates why accusative pronouns are harder. The patterns we demonstrate (both for performance and bias) could arise from a number of sources beyond mere frequency, including quirks of our dataset, or the distribution of semantic roles in training data for coreference resolution systems.
Second, we echo prior calls for fairness researchers to attend to the differences between social gender and terms that index it (Cao and Daumé III, 2021; Gautam et al., 2024b), to include more diversity in pronouns (Baumler and Rudinger, 2022; Lauscher et al., 2022; Hossain et al., 2023), and to move towards richer operationalizations of gender (Devinney et al., 2022; Ovalle et al., 2023) and bias (Blodgett et al., 2020). Specifically, future work on bias in coreference resolution should treat pronominal bias as distinct from (social) gender bias, defend how and why pronouns are mapped to social gender, and move beyond binary, oppositional methods of evaluation.
Lastly, as our work is a case study in how careful data curation and operationalization affects claims about system performance and bias, we emphasize the need for thoughtful data work (Sambasivan et al., 2021), and encourage the use of automatic checks when feasible, as in our work.
7 Related work
Our work is also one among several papers that investigate datasets for problems including low quality or noisy data (Elazar et al., 2024; Abela et al., 2024), artifacts (Shwartz et al., 2020; Herlihy and Rudinger, 2021; Elazar et al., 2021; Dutta Chowdhury et al., 2022), contamination Balloccu et al. (2024); Deng et al. (2024), and issues with conceptualization and operationalization of bias (Blodgett et al., 2021; Selvam et al., 2023; Nighojkar et al., 2023; Subramonian et al., 2023; Gautam et al., 2024b). We cover many of these areas, but do not control for dataset artifacts, which we explain further in our Limitations section.
@@ 8 Conclusion
We demonstrate a number of issues with the well-known Winogender Schemas dataset, which we fix in our new, expanded WinoPron dataset. In addition, we propose a novel way to evaluate pronominal bias in coreference resolution that goes beyond the binary and focuses on lexical gender. With our new dataset, we evaluate both supervised coreference resolution systems and language models, and find that the grammatical case of pronouns affects model performance and bias, and that bias varies widely across models, pronoun sets and grammatical cases. Our work demonstrates that measurements of bias and robustness are only as good as the datasets and metrics we use to measure them, and we call for careful attention when developing future resources for evaluating bias and coreference resolution, with attention to grammatical case, more careful operationalizations of bias, and greater diversity in the pronouns we consider.
@@
As in Winogender Schemas, our schemas are not “Google-proof” and could conceivably be solved with heuristics, including word co-occurrences, which is a primary concern when creating and evaluating Winograd schemas (Levesque et al., 2012; Amsili and Seminck, 2017; Elazar et al., 2021). The fact that we do not control for this means that our dataset gives generous estimates of system performance, particularly for strong language models like FLAN-T5, but it also means that this dataset is inappropriate to test “reasoning.” Our dataset construction instead controls for simple system heuristics that are relevant for coreference resolution, such as always picking the first entity in the sentence, or always picking the second.
We take steps to prevent data contamination (Jacovi et al., 2023), including not releasing our data in plain text, and not evaluating with language models behind closed APIs that do not guarantee that our data will not be used to train future models (Balloccu et al., 2024). However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on WinoPron with models trained on data after August 2024.
Finally, we note that as our evaluation set only contains one set of templates per occupation-participant pair, our results represent a point in the distribution of bias related to that occupation. We thus echo Rudinger et al.’s (2018) view of Winogender Schemas as having “high positive predictive value and low negative predictive value” for bias. In other words, they may demonstrate evidence of pronominal bias in systems, but not prove its absence. In the case of large language models in particular, using a small number of templates for templatic evaluation is known to be brittle even to small, meaning-preserving changes to the template (Seshadri et al., 2022; Selvam et al., 2023). Our dataset’s small size is a result of us requiring a tightly controlled and structured dataset to evaluate how coreference resolution varies. Thus, it may differ from realistic examples (which would have other differences that confound bias results). We wish to emphasize that in addition to controlled datasets like ours, realistic evaluation is also necessary for holistically evaluating performance, robustness and bias in coreference resolution.
@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, and are grateful to Rachel Rudinger, Benjamin Van Durme, and our CRAC reviewers for their comments. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Anne Lauscher’s work is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A List of Occupations
@@
Appendix B Annotator Demographics
@@
Appendix C Annotation Instructions
C.1 Task 1 Description
@@ C.1.1 Grammaticality
C.1.2 Questions about pronouns
@@ C.2 Task 2 Description
Appendix D Prompting
@@
Table 7: Prompting templates, where “task” is filled with each dataset instance, “pronoun” is the unique third person singular pronoun in that dataset instance, and “options” are the occupation and the participant.
  1. 1.
    No options
  2. 2.
    The occupation is presented first and the participant second
  3. 3.
    The participant is presented first and the occupation second
Appendix E Additional Results
We report additional results on double- and single-entity sentences in WinoPron: $F_{1}$ scores in Table 8, precision in Table 9, and recall in Table 10. Note that FLAN-T5 models generally perform worse on single-entity sentences compared to double-entity sentences because some of our prompts include options (see Section D for details) that confuse the model in this setting, despite being necessary to resolve double-entity sentences.
@@ Appendix E Additional Results
Double-entity sentences
All 64.4 67.3 61.6 70.1 51.6 62.4 78.0 89.0 88.8
Nominative 73.5 77.6 67.2 77.2 51.9 65.4 85.1 94.7 96.7
Accusative 52.2 57.5 54.6 59.5 50.4 58.4 69.9 82.5 79.1
Possessive 67.4 66.5 62.9 73.6 52.3 63.4 79.1 89.7 90.7
he/him/his 79.2 79.6 62.8 71.5 51.5 64.1 81.5 88.8 90.2
she/her/her 76.3 76.6 62.1 71.6 51.5 66.1 83.3 90.6 89.9
they/them/their 67.5 63.7 61.2 68.9 51.8 60.5 77.0 88.6 88.0
xe/xem/xyr 8.5 38.6 60.4 68.5 51.4 58.7 70.3 88.0 87.3
Single-entity sentences
All 73.2 75.6 95.5 88.0 77.3 76.3 81.5 83.1 84.3
Nominative 80.0 82.5 99.5 99.3 78.3 80.8 89.8 93.3 97.0
Accusative 61.1 65.0 87.3 67.5 76.2 69.6 69.8 70.1 66.5
Possessive 77.1 78.0 99.8 97.1 77.5 78.5 84.7 85.7 89.2
he/him/his 92.7 94.3 94.7 85.6 77.6 81.3 86.8 88.2 88.6
she/her/her 90.9 91.6 96.2 88.9 77.4 81.1 87.6 88.8 87.1
they/them/their 75.2 69.8 96.0 88.7 79.3 76.1 84.3 85.7 86.8
xe/xem/xyr 2.2 27.3 95.2 88.7 75.0 66.3 67.0 69.4 74.6
Table 8: F1 of coreference resolution systems on double- and single-entity sentences in WinoPron. We report F1 overall, and split by grammatical case and pronoun set. Red, italicized numbers are worse than chance (50.0 for double-entity sentences and not applicable for single-entity sentences).
Data LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
Double-entity sentences
All 79.1 80.1 62.1 70.6 51.9 62.9 78.4 89.5 89.4
Nominative 88.3 88.7 67.4 77.4 52.1 65.7 85.4 95.1 97.1
Accusative 63.4 67.9 55.3 59.9 50.7 58.8 70.2 83.2 79.5
Possessive 86.1 83.6 63.5 74.3 52.8 64.3 79.6 90.2 91.5
he/him/his 79.7 80.1 63.0 71.6 51.7 64.3 81.8 89.3 90.6
she/her/her 77.6 77.9 62.3 71.8 51.7 66.3 83.6 91.1 90.3
they/them/their 79.1 80.2 61.8 69.5 52.0 60.8 77.3 89.0 88.5
xe/xem/xyr 100.0 88.1 61.3 69.3 52.1 60.1 70.7 88.6 88.0
Single-entity sentences
All 100.0 100.0 96.0 88.4 78.9 77.6 82.4 84.0 85.6
Nominative 100.0 100.0 100.0 100.0 79.3 81.6 90.4 93.9 97.4
Accusative 100.0 100.0 88.1 67.9 77.5 70.5 70.7 71.1 68.1
Possessive 100.0 100.0 99.8 97.1 79.8 80.7 85.9 86.8 90.8
he/him/his 100.0 100.0 95.0 86.0 78.6 81.9 87.5 88.9 89.5
she/her/her 100.0 100.0 96.4 89.1 78.5 81.7 88.1 89.4 87.9
they/them/their 100.0 100.0 96.4 89.1 80.3 76.9 85.2 86.5 87.9
xe/xem/xyr 100.0 100.0 96.3 89.3 77.9 69.2 68.3 70.9 76.7
Table 9: Precision on double- and single-entity sentences overall, and split by grammatical case and pronoun set. Red, italicized numbers are worse than chance (50.0 for double-entity sentences, N/A for single-entity sentences).
Data LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
Double-entity sentences
All 54.2 58.0 61.1 69.7 51.3 61.9 77.7 88.5 88.3
Nominative 62.9 69.0 67.1 77.1 51.8 65.2 84.8 94.3 96.3
Accusative 44.4 49.8 54.0 59.2 50.2 58.0 69.6 81.9 78.6
Possessive 55.4 55.2 62.3 72.9 51.9 62.5 78.7 89.3 90.0
he/him/his 78.6 79.2 62.5 71.4 51.4 63.9 81.1 88.3 89.7
she/her/her 75.0 75.3 61.9 71.4 51.4 65.9 83.0 90.1 89.5
they/them/their 58.9 52.8 60.6 68.3 51.6 60.2 76.8 88.1 87.5
xe/xem/xyr 4.4 24.7 59.4 67.8 50.8 57.4 69.9 87.5 86.6
Single-entity sentences
All 57.8 60.8 95.1 87.6 75.9 75.0 80.6 82.1 83.1
Nominative 66.7 70.2 99.0 98.5 77.3 80.0 89.2 92.7 96.6
Accusative 44.0 48.1 86.5 67.1 74.9 68.7 69.0 69.1 65.0
Possessive 62.7 64.0 99.8 97.1 75.4 76.4 83.5 84.6 87.6
he/him/his 86.4 89.2 94.4 85.3 76.5 80.8 86.1 87.4 87.8
she/her/her 83.3 84.4 96.1 88.6 76.2 80.5 87.1 88.2 86.2
they/them/their 60.3 53.6 95.6 88.3 78.3 75.3 83.5 85.0 85.8
xe/xem/xyr 1.1 15.8 94.2 88.1 72.4 63.7 65.7 68.0 72.5
Table 10: Recall on double- and single-entity sentences overall, and split by grammatical case and pronoun set. Red, italicized numbers are worse than chance (50.0 for double-entity sentences, N/A for single-entity sentences)